OpenStudy (anonymous):

Identify the missing factors: -14x = (7)( )

OpenStudy (anonymous):

-2x

OpenStudy (anonymous):

same thing as the other topic. Prented the missing factor =y -14x = (7) (y) y= -14x/7 y = -2x